OpenStudy (anonymous):

Find the partial sum for a_n = 2/(3^n)

OpenStudy (anonymous):

Partial sum for the nth term

OpenStudy (anonymous):

its a g.p with a=2,r=1/3

jhonyy9 (jhonyy9):

- partial sum for n= ? a_1=2/3 a_2=2/9 - a_1 +a_2 =2/3 +2/9 =(6+2)/9 =8/9 a_3=2/27 - 8/9 +a_3 =8/9 +2/27 =(24+2)/27 =26/27 a_4=2/81 - 26/27 +a_4 =26/27 +2/81 =(78+2)/81 =80/81 - - - a_n=2/(3^n) -------------- - so than from these result that a_1 + ... +a_n = ((3^n) -1)/(3^n) - hope that is understandably - good luck - bye !

jhonyy9 (jhonyy9):

how do you like it ?

OpenStudy (anonymous):

If you start with n=1 \[s_k=3^{-k} \left(3^k-1\right) \] If you start with n=0 \[ s_k=3^{-k} \left(3^{k+1}-1\right) \]